# Faculty of Computer Science, TU Dresden, Germany

### v2

@@ 4 The Riskman Ontology & Shapes: Overview and Usage
The Riskman ontology was developed within a consortium of domain experts in medical devices and risk management, ontology engineers and software developers. For the development of the ontology, we utilised the Linked Open Terms (LOT) methodology [58], a framework for ontology development that builds upon the NeOn methodology [65]. It comprised the following activities:
@@
Requirement specification. We studied relevant norms/standards and inspected risk management files of real medical devices to gain a shared understanding of the required terms. This was accompanied by several meetings within the consortium and workshops with domain experts from the field: risk managers, manufacturers, consultants, and notified bodies. As a result, objectives and a glossary of terms were compiled, forming the basis for the conceptual ontology model. Moreover, the VDE Spec 90025 [7] working group was formed to develop a specific submission format and its requirements based on the conceptual model.
@@
Implementation. We encoded the ontology and shapes using OWL and SHACL. During development, we used the HermiT reasoner [63] and the OntOlogy Pitfalls Scanner (OOPS!) [59] to ensure the ontology’s validity and consistency.
@@
Publication and Maintenance. The Riskman ontology, shapes and documentation are published online with a permanent URL provided by w3id. The ontology is maintained in a GitHub repository [8] for issue tracking and version control.
@@
As mentioned before, the approach assumes risk documentation to be submitted as HTML files with data encoded in RDFa.
@@
It is then extracted using an RDF distiller and used as input to an materialisation reasoner alongside with the Riskman- and potentially additional ontologies.
@@
an enhanced DL knowledge base is conceived, which is later validated for conformity against the Riskman SHACL constraints (and potentially additional constraints) utilising a SHACL validator.
@@
We provide a prototypical implementation of the validation pipeline together with examples showcasing validation results [30].
@@
In the following subsections, we will explain the approach in more detail and provide preliminaries necessary for understanding. Afterwards, the actual Riskman ontology and shapes will be shown. This section will conclude with showcasing the approach on our running GIIP example and giving hints of how the approach can be extended depending on further needs of interested parties.
@@ 4.1 The Description Logic
Firstly, it is important to note that all of the Riskman ontology can be represented within the description logic [10, 11]. We therefore briefly recall its syntax, semantics, and aspects of reasoning that are relevant to our approach.
@@
As usual for description logics, the semantics of is defined via interpretations with a non-empty domain and an interpretation function . An interpretation is a model of an ABox (TBox ) iff it satisfies all elements of () as per Table 2. An assertion is entailed by , written , iff every model of is a model of . We also remark that RIAs (13)–(16) satisfy the syntactic restriction imposed by Baader et al. [11, Section 3].
There are two kinds of reasoners, predominantly ontology classifiers [43]. The first kind implements tableau calculi while the second is based on materialisation. Tools like ELK [43], also distributed with ontology modelling frontends like Protégé [53], follow the materialisation-based approach by step-wise computing all logical consequences from a given ABox (i.e., the risk management information) and an ontology (i.e., Riskman). The polynomial-time complexity of reasoning [10, 11] opens the stage for general purpose Datalog reasoners (e.g. nemo [41]) to implement materialisation [15]. Thus, a wide variety of highly optimised tools are at the disposal of potential users of Riskman.
4.2 Classes and Properties
Riskman’s classes and their interrelationships are depicted in the schema diagram in Fig. 2 (top), where also domains and ranges of properties can be read off the edges. The main axioms of the Riskman ontology are given in Fig. 2 (bottom), including detailed superclass declarations, and relationships between properties via role inclusion axioms, role chains, or transitivity statements.
4.3 Ontology Design Patterns
We opted for a lightweight ontology that captures the outcomes of performed risk management with low ontological commitment, as the particular needs might differ from manufacturer to manufacturer. To ensure the usability and extensibility of the ontology, we employed the Stub Metapattern [48, 49], which “acts as a type of placeholder for future extensions.” Specifically, the class DeviceProblem is intended to link to the IMDRF’s controlled vocabularies [39, Annex A]; another stub, HazardousSituation, enables to potentially reuse the Hazardous Situation Pattern [50, 16] should a more fine-grained representation be required.
@@ 4.4 Probabilities and Severities
VDE Spec 90025 (importing from ISO 14971) defines risk as the “combination of the probability of occurrence of harm and the severity of that harm”, and so to represent concrete risks it is necessary to also represent concrete “values” for probability (and severity). Since mathematical probabilities are virtually impossible to obtain for events and situations that are hypothetical from the outset, the typical approach in risk management (cf. Sect. 3) is to use a fixed, finite number of probability magnitudes, each representing an interval of real-valued probabilities, and being naturally ordered on a logarithmic scale, e.g. “improbable”: , “remote”: , “occasional”: , “probable”: , and “frequent”: . The exact number of different magnitudes varies and is up to the manufacturer, but choosing (as above) is common; a similar approach is typically also used to represent severity.
@@
This introduces not only the discretised probability values (e.g. for we would get that and ), but also an ordering on these values. Most importantly, however, the GCIs introduced above implement the “computation” of overall probability from probabilities and [40]. Essentially, the multiplication works by adding exponents of upper bounds of intervals, e.g. results in , and all possible computations for the given (and ) are expressed via GCIs.
@@ 4.5 Shapes
As explained above, our ontology can be used to derive implicit information via reasoning. The materialised graph is then stored in RDF format [72]. (As an illustration, we depict the implementation of Example 1 in Fig. 4.) For our use case, the evaluation of risk reports, it is now important that we are also able to check for missing information (risks could be not mitigated) or for mismatches between values (a mitigation should not increase severity or probability of a risk). These kinds of checks on a (hopefully) complete and self-contained risk report are conceptually not a good fit with OWL’s open-world assumption. We thus define SHACL shapes that operate on the materialised RDF graph and implement the most important requirements risk reports should fulfil. We describe them in what follows. Users clearly can, and most likely will, add custom shapes.
We adopt the abstract syntax of SHACL constraints proposed by Corman et al. [18, 6], capturing the core components of the SHACL specification [71]. For our purposes, we conveniently re-use description logic vocabulary, viz., pairwise disjoint sets of classes, of properties, and of individuals. A finite set of assertions (an ABox) can then be seen as representing a labelled graph, with individuals acting as nodes, classes labelling nodes, and properties labelling edges. The syntax of shape expressions and path expressions is shown in Fig. 3 (top). For the semantics, a given graph (ABox) with nodes (individuals) defines an evaluation function that assigns to each path expression a binary relation , and to each shape expression a set via structural induction as shown in Fig. 3 (middle).
A shape constraint is an expression of the form , with and a shape expression. A shape schema is a pair where is a set of shape constraints and is a set of target concept assertions. Intuitively, a target expresses the requirement that be labelled by . Formally, an ABox is a model for a set of constraints iff for all . An ABox is validated against a schema iff there exists a set of concept assertions such that (1) , (2) , and (3) is a model for .
The shapes constraints that complement the Riskman ontology are listed in Fig. 3 (bottom). To illustrate their use, in Fig. 4 individuals satisfying the shape expression part of a constraint are labelled by the constraint’s number.
@@
4.6 Example and Extensibility
The example from Fig. 4 illustrates how our ontology and shapes could be used in practice. Alas, even a carefully designed framework like ours, which takes the needs of the different stakeholders into account, can never meet everyone’s wishes. Companies could have individual requirements, authorities might want to add specialised tests, and regulations differ per region and change over time. This is one of the main reasons why we chose Semantic Web technologies for our task: they are easy to extend and can be combined with existing data. To illustrate the extensibility of our approach, we give a small example: We described earlier (Sect. 4.4) that companies often define magnitude levels to model probabilities and severities of risks. While studying risk reports, we observed that these definitions sometimes come with a so-called risk acceptance matrix, a schema indicating which combinations of probability and severity a company considers critical. Such information can be modelled with our ontology by adding a class . Assume for the sake of example that the combination “probability and severity ” is critical. This can be indicated by an axiom
allowing to easily check for controlled risks with critical risk levels using SHACL:
Of course, more complex additions and constraints are also possible.
@@ 5 Discussion and Outlook
We presented the Riskman ontology & shapes with their intended use of representing and analysing risk management information for medical devices. Analysed risks and their mitigations are represented as an ABox, the Riskman ontology is used with a materialisation engine to infer implicit knowledge, and lastly SHACL constraints are used to check whether the input data conforms to given requirements. The resource (ontology + shapes constraints) is available, including a reference implementation of the whole pipeline. With feedback from manufacturers and notified bodies incorporated into it, we envision Riskman to improve the work lives of risk managers and certification auditors alike.
@@
The issue of medical device safety will only become more important in the future, especially with further digitisation [66]. An interesting next step for the Riskman ontology is to not only assess submission completeness, but also evaluate the quality of the assurance [17]. First steps, including a proof of concept, have recently been achieved in the context of assurance cases [28], but are known to be hard to generalise [21]. Possible techniques to consider for evaluation of risk management artifacts represented using Riskman are dialogue-based approaches to proof theory in structured argumentation [24], as well as other approaches to provide justifications in logic-based knowledge representation formalisms [23].
In the future, we also want to use existing data about adverse events (partially annotated with terms from IMDRF taxonomies, e.g. from the FDA’s Manufacturer and User Device Facility Experience database, MAUDE [70]) to allow for a more informed evaluation of “whether all hazards have been identified“, a classical point of risk management that cannot strictly be proved by a manufacturer, but potentially be disproved by a sufficiently knowledgeable reviewer/auditor.
@@
Ahmetaj et al. [3] analysed how non-validation of SHACL constraints can be explained to users (in terms of repairs), which can potentially be applied to our work and be included in a future Riskman-based work bench for risk managers.
@@
Instead of delegating the inference and validation steps to two different services (reasoners/validators), we could also utilise advancements in the area of combining OWL+SHACL, e.g. by converting the ontology and shapes into a single set of SHACL constraints [4]. While this involves an exponential blowup in general [4], the fact that Riskman stays within might constitute an interesting special case. Alternatively, having a single reasoner perform inferencing as well as constraint checking could also be achieved by translating ontology axioms and shapes constraints into answer set programming [14]. With different implementations in place, we will experimentally evaluate and compare them.
@@
As a possible alternative to extending some concepts that are currently stubs (in the sense of the stub metapattern [48]), we envision to use the novel formalism of standpoint logic [33, 34] to import and attach further ontologies to Riskman. In this regard it is especially notable and useful that the combination of standpoint logic and the description logic retains the latter’s polytime computational complexity [36, 35]. A prominent candidate for integration is the US National Cancer Institute’s thesaurus (NCIt) [54], which can be expressed in [11].
@@ Acknowledgements.
This work was supported by funding from BMBF within projects KIMEDS (grant no. GW0552B), MEDGE (grant no. 16ME0529), and SEMECO (grant no. 03ZU1210B).
Appendix 0.A Definitions of Terms from VDE Spec 90025
In this section, we briefly recall relevant definitions from VDE Spec 90025 [7]. Table 3 provides definitions of terms borrowed from ISO 14971 [40], while table 4 gives definitions of terms newly introduced in VDE Spec 90025.

### v3

@@ 4 The Riskman Ontology and Shapes: Overview and Usage
The Riskman ontology was developed taking into account requirements and conceptual design choices laid out by a consortium of domain experts in medical devices and risk management, ontology engineers, and software developers. For ontology development, we used the Linked Open Terms (LOT) methodology [Poveda-Villalon2022], an ontology development framework that builds upon the NeOn methodology [SurezFigueroa2015TheNM]. It comprised the following:
@@
Requirement specification. Studying relevant norms/standards and inspecting real medical devices’ risk management files to gain a shared understanding of the required terms. This was accompanied by several meetings within the consortium and workshops with domain experts from the field: risk managers, manufacturers, consultants, and notified bodies. As a result, we compiled objectives and a glossary of terms, forming the basis for the conceptual ontology model. In parallel, the VDE Spec 90025 [VDESpec24] working group developed a technical submission format with constraints based on the conceptual model.
@@
Implementation. We encoded the ontology and shapes using OWL and SHACL. During development, we used the HermiT reasoner [shearer2008hermit] and the OntOlogy Pitfalls Scanner (OOPS!) [poveda2014oops] to ensure the ontology’s validity and consistency.
@@
Publication and Maintenance. The Riskman ontology, shapes, and documentation are published online with a permanent URL provided by w3id. To assess the conformity of the ontology to the FAIR (Findable, Accessible, Interoperable, and Reusable) principles [wilkonson2016fair], we used the Ontology Pitfall Scanner for the FAIR Principles (FOOPS!) [foops2021]. Riskman is maintained in a GitHub repository [RiskmanRepo] for issue tracking/version control.
@@
As previously noted, the approach expects risk documentation to be submitted as HTML files containing risk management data encoded as RDF triples.
@@
An RDF distiller extracts the RDF graph from the HTML by removing the markup and isolating the RDF triples, which are then used as input to an reasoner alongside Riskman (and potentially additional ontologies).
@@
a materialized DL knowledge base is obtained, which is then validated against the Riskman SHACL constraints (and potentially additional constraints) utilizing a SHACL validator.
@@
We provide a prototypical implementation of the validation pipeline together with examples showcasing validation results [RiskmanPipeline].
@@
In the following subsections, we explain the approach in more detail and provide preliminaries necessary for understanding. Afterwards, we present the actual Riskman ontology and shapes. The section concludes with illustrating the approach on our running example and showing how it can be extended to further needs of interested parties.
@@ 4.1 The Description Logic
Firstly, it is important to note that all the Riskman ontology can be represented within the description logic [BaaderBL05, BaaderLB08]. We therefore briefly recall its syntax, semantics, and aspects of reasoning that are relevant to our approach.
@@
As usual for description logics, the semantics of is defined via interpretations with a non-empty domain and an interpretation function . An interpretation is a model of an ABox (TBox ) if it satisfies all elements of () as per Table 2. An assertion is entailed by , written , if every model of is a model of . We also remark that RIAs (12)–(15) satisfy the syntactic restriction imposed by Baader et al. [BaaderLB08, Section 3].
There are two kinds of reasoners, predominantly ontology classifiers [KKS:ELK2013]. The first kind implements tableau calculi while the second is based on materialisation. Tools like ELK [KKS:ELK2013], also distributed with ontology modelling frontends like Protégé [Musen15], follow the materialisation-based approach by step-wise computing the relevant logical consequences of a given ABox and an ontology. In the context of Riskman, the “relevant” consequences are all concept and role assertions involving concept, role, and individual names occurring in the input ABox, where concept assertions must be atomic, that is, of the form with . The fact that no new individuals are required for this process is guaranteed by Riskman’s design, which employs existential quantification only on the left-hand side of general concept inclusion axioms.111The optional add-on of Section 4.3 has existential restriction on right-hand sides, but only involving nominal concepts, thus no new individuals need to be created during inference. The completion (addition of all relevant consequences) of a given ABox w.r.t. the Riskman ontology can therefore (also due to the satisfaction of the syntactic restriction by Baader et al. [BaaderLB08, Section 3]) be computed in polynomial time [BaaderBL05, BaaderLB08], which enables general purpose Datalog reasoners (e.g. Nemo [nemo24]) to implement Riskman ABox completion [CDK20:DatalogEL]. Thus overall, a variety of optimized tools are at the disposal of potential Riskman users.
4.2 Classes, Properties, and Ontology Design Patterns
The main axioms of the Riskman ontology are given in Figure 2, including detailed superclass declarations, and relationships between properties via role inclusion axioms. Riskman’s classes and their general interrelationships are depicted in the schema diagram in Figure 3, where also domains and ranges of properties can be read off the edges.
We opted for a lightweight ontology that captures the outcomes of risk management with low ontological commitment, as the particular needs might differ from one manufacturer to another. To ensure the usability and extensibility of the ontology, we employed the Stub Metapattern [KrisnadhiH16, StubURL], which “acts as a type of placeholder for future extensions.” Specifically, the class DeviceProblem is intended to link to the IMDRF’s controlled vocabularies [IMDRF20AET, Annex A]; another stub, HazardousSituation, enables to reuse the Hazardous Situation Pattern [Lawrynowicz2015TheHS, Cheatham2016AMT] if a more fine-grained representation is required.
@@ 4.3 Probabilities and Severities
VDE Spec 90025 (importing from ISO 14971) defines risk as the “combination of the probability of occurrence of harm and the severity of that harm”, and so to represent concrete risks it is necessary to also represent concrete “values” for probability (and severity). Mathematical probabilities are virtually impossible to accurately determine for events and situations that are hypothetical from the outset. Therefore, the typical approach in risk management (cf. Section 3) is to use a fixed, finite number of probability magnitudes, each representing an interval of real-valued probabilities, and being naturally ordered on a logarithmic scale, e.g. “improbable”: , “remote”: , “occasional”: , “probable”: , and “frequent”: . The exact number of different magnitudes varies and is up to the manufacturer; choosing (as above) is common. A similar approach is typically also used for severity.
@@
This introduces not only the discretized probability values (e.g. for we get and ), but also an ordering on these values. Most importantly, however, the GCIs in implement the “computation” of overall probability from probabilities and [ISO14971]. Essentially, the multiplication works by adding exponents of upper bounds of intervals, e.g. yields , and all possible computations for the given are expressed via GCIs.
@@ 4.4 Shapes
As explained above, our ontology can be used to derive implicit information via reasoning. The materialized graph is then stored in RDF format [owlov]. (As an illustration, we depict the implementation of 1 in Figure 5.) Our use case, the evaluation of risk reports, requires the capability to check for missing information or for mismatches between values (e.g., whether a mitigation of a risk does not increase its severity or probability). These kinds of checks on a (hopefully) complete and self-contained risk report are conceptually not a good fit with OWL’s open-world assumption. We thus define SHACL shapes that operate on the materialized RDF graph and implement the most important requirements risk reports should fulfil. We describe them in what follows.
We adopt the abstract syntax of SHACL constraints proposed by Corman et al. [CormanRS18, AndreselCORSS20], capturing the core components of the SHACL specification [SHACL]. For our purposes, we conveniently re-use description logic vocabulary, viz., pairwise disjoint sets of classes, of properties, and of individuals. A finite set of assertions (an ABox) can then be seen as representing a labelled graph, with individuals acting as nodes, classes labelling nodes, and properties labelling edges. The syntax of shape expressions and path expressions is shown in Table 3 (top). For the semantics, a given graph (ABox) with nodes (individuals) defines an evaluation function that assigns to each path expression a binary relation , and to each shape expression a set via induction as shown in Table 3 (bottom).
@@
A shape constraint is an expression of the form , with and a shape expression. A shape schema is a pair where is a set of shape constraints and is a set of target concept assertions. Intuitively, a target expresses the requirement that be labelled by . Formally, an ABox is a model for a set of constraints if for all . An ABox is validated against a schema if there exists a set of concept assertions such that (1) , (2) , and (3) is a model for . The shapes constraints that complement the Riskman ontology are listed in Figure 4. To illustrate their use, in Figure 5 individuals satisfying the shape expression part of a constraint are labelled by the constraint’s number.
Constraint 19 encodes checks for non-increasing residual risk levels, i.e. whether the probability or severity after implementing a mitigation is not higher than before. Constraints 16, 18, 20, and 21 are duals of the subclass declarations for , , , and , respectively, in that they enforce instances of these classes to contain all the necessary components of the class definitions, as required by VDE Spec 90025. Constraint 17 encodes that (i) every sub- of an must be an and (ii) an must have a . Assuming the ABox has previously been materialized not only by means of the Riskman ontology, but also the additional “probability-severity” ontology , Constraint 21 requires that (i) an overall probability is present (specified directly or inferred from probabilities and via ), (ii) that holds in case all three have been specified, and (iii) a severity magnitude is present. Finally, the check whether all leaf nodes of the SDA tree are s is provided by Constraint 22. Figure 5 illustrates how to use our ontology and shapes by implementing the example of Figure 1.
4.5 Extensibility
Alas, even a carefully designed framework like ours, which takes the needs of the different stakeholders into account, can never meet everyone’s wishes. Companies could have individual requirements, authorities might want to add specialized tests, and regulations differ per region and change over time. To illustrate the extensibility of our approach, we give an example: We described in Section 4.3 that companies often define magnitude levels to model probabilities and severities of risks. In practice, these definitions sometimes come with a so-called risk acceptance matrix, a schema indicating which combinations of probability and severity a company considers critical. This can be modelled with our ontology by adding a class . Assume now for the sake of example that combination “probability and severity ” is critical. This can be stated via an axiom allowing to easily check for controlled risks with critical residual risk levels using SHACL via a constraint . Of course, more complex additions of further shapes and constraints are likewise possible.
@@ 5 Discussion and Outlook
We presented the Riskman ontology and shapes with their intended use of representing and analysing risk management information for medical devices. Analyzed risks and their mitigations are represented as an ABox, the Riskman ontology is used with a reasoner to infer implicit knowledge, and lastly SHACL constraints are used to check whether the input data conform to given requirements. The ontology and shapes constraints are freely available, including a reference implementation of the whole pipeline. With feedback from manufacturers and notified bodies incorporated into it, we envision Riskman to improve the work lives of risk managers and certification auditors alike.
@@
The issue of medical device safety will only become more important in the future, especially with further digitization [SujanSC19]. This is the case even more so for devices that use artificial intelligence themselves, and as such fall under the regulations of the AI Act [AIAct]; an integration with the like-minded AIRO ontology [GolpayeganiPL22] is an important topic for future work. An interesting next step for the Riskman ontology is to not only assess submission completeness, but also evaluate the quality of the assurance [ChowdhuryWPL20]. First steps, including a proof of concept, have recently been achieved in the context of assurance cases [FosterNGWK21], but are known to be hard to generalize [delaVaraJMP19]. Possible techniques to consider for evaluation of risk management artifacts represented using Riskman are dialogue-based approaches to proof theory in structured argumentation [DillerGG21], as well as other approaches to provide justifications in logic-based knowledge representation formalisms [DeneckerBS15].
@@
Ahmetaj et al. [AhmetajDOPSS21] analyzed how non-validation of SHACL constraints can be explained to users (in terms of repairs), which can potentially be applied to our work and be included in a future Riskman-based work bench for risk managers.
@@
Instead of delegating the inference and validation steps to two different services (reasoners/validators), we could also utilize advancements in the area of combining OWL and SHACL, e.g. by converting the ontology and shapes into a single set of SHACL constraints [AhmetajOOS23]. While this involves an exponential blowup in general [AhmetajOOS23], the fact that Riskman stays within might constitute an interesting special case. Alternatively, having a single reasoner perform inferencing as well as constraint checking could also be achieved by translating ontology axioms and shapes constraints into answer set programming [BrewkaET11]. Different implementations could then be compared experimentally.
@@
As a possible alternative to extending some concepts that are currently stubs (in the sense of the stub metapattern [KrisnadhiH16]), we envision to use the novel formalism of standpoint logic [AlvarezR21, AlvarezRS22] to import and attach further ontologies to Riskman. In this regard it is especially notable and useful that the combination of standpoint logic and the description logic retains the latter’s polynomial time computational complexity [AlvarezRS23a, AlvarezRS23b]. A prominent candidate for integration is the US National Cancer Institute’s thesaurus (NCIt) [ncitURL], which can be expressed in [BaaderLB08]. On the other hand, Riskman itself could also be embedded into a top-level ontology, e.g. BFO [Arp15] or GFO [Herre10].
@@ Acknowledgements.
This work was supported by funding from BMFTR (Federal Ministry of Research, Technology and Space) within projects KIMEDS (grant no. GW0552B), MEDGE (grant no. 16ME0529), SEMECO (grant no. 03ZU1210B), and SECAI (within DAAD project 57616814, School of Embedded Composite AI, https://secai.org/, as part of the program Konrad Zuse Schools of Excellence in Artificial Intelligence).
We are indebted to numerous people who explained the theory and practice of risk management, commented on our earlier formalisations, and participated in discussions that improved our understanding of the domain. In alphabetical order, we thank: Philipp Bank, Manuel Baur, Samet Bayraktar, Gustav Bieberstein, Nidhal Chouchane, Tony Dietrich, Anne Eßlinger, Ludger Evers, Felix Gebhardt, Stephen Gilbert, Andreas Halbleib, Christian Helmbold, Evi Hartig, Aleksandr Ilinykh, Abtin Jamshidirad, Simon Kilcher, Jürgen Koch, Tina Küttner, Andreas Lämmerzahl, Felix Lempke, Alf Ludwig, Bokhodir Mamadaliev, Svetlana Miasoedova, Davood Moghadas, Martin Neumann, Andreas Purde, Katharina Rehde, Christian Rosenzweig, Kerstin Rothweiler, Jasmine Schirmer, Hubert Sefkovicz, Sebastian Schostek, Robin Seidel, Robert Stelzmann, Sarah Tsurkan, Heike Vocke, Hans Wenner, Martin Witte, Sven Wittorf, Juliane Wober, and Uwe Zeller.
Appendix A Definitions of Terms from VDE Spec 90025
In this section, we briefly recall relevant definitions from VDE Spec 90025 [VDESpec24]. Table 4 provides definitions of terms borrowed from ISO 14971 [ISO14971], while table 5 gives definitions of terms newly introduced in VDE Spec 90025.